OpenStudy (anonymous):

Find two numbers with a sum of 20 and a difference of 14. A. 4 and 18 B. –3 and –17 C. 2 and 16 D. 3 and 17

Parth (parthkohli):

\( \color{Black}{\Rightarrow a + b = 20}\) \( \color{Black}{\Rightarrow a - b = 14}\)

Parth (parthkohli):

Solve this system.

Parth (parthkohli):

Try solving it by elimination

OpenStudy (anonymous):

Why dont you just and try to use the solutions and see which yields true? This just comes accross as lazy...

OpenStudy (anonymous):

Use simultaneous equations: a+b=20 a-b=14 take them away from eachother: 2b=6 b=3 substitute into one of the equations: a+3=20 a=17

OpenStudy (anonymous):

From the sum being 20 alone you can tell the answer is D. Literally zero effort and all I did was sum two numbers for each possibility.

Parth (parthkohli):

The main thing is learning the algebraic solutions. That is better.

OpenStudy (anonymous):

the answer is D

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

nor problem :)